# Interpolating neural network: A lightweight yet precise architecture for data training, equation solving, and parameter calibration

### v4

@@ Abstract
Artificial intelligence (AI) has revolutionized software development, shifting from task-specific codes (Software 1.0) to neural network-based approaches (Software 2.0). However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations. We introduce an interpolating neural network (INN), grounded in interpolation theory and tensor decomposition, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration. INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Demonstrated in metal additive manufacturing, INN rapidly constructs an accurate surrogate model of Laser Powder Bed Fusion (L-PBF) heat transfer simulation, achieving sub-10-micrometer resolution for a 10 mm path in under 15 minutes on a single GPU. This makes a transformative step forward across all domains essential to engineering software.
@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 partially resolved the issue of labor-intensive programming in Software 1.0 and has significantly advanced the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies in the fields of engineering and science presents unique challenges. One such challenge is the relatively low generalization accuracy of machine learning (ML)-based solvers compared to traditional numerical methods [4, 5, 6, 7, 8]. After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9]. Another challenge is the curse of dimensionality that arises when solving inverse design problems. As the design parameter space grows, computation requirements easily become prohibitive. Furthermore, the complexity of physical simulations continues to increase. For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis within a reasonable simulation time [10, 11, 12].
@@
In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of computational and data-driven engineering software into a new paradigm.
@@
“Engineering Software 2.0 is an end-to-end software system that unifies training, solving, and calibrating in science/engineering problems."
@@
We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0. The key ideas behind INNs are: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function. Figure 1 shows a pictorial description of the approach. When the input domain is discretized with a regular mesh (see the special case in Figure 1), an INN leverages tensor decomposition (TD) [13, 14, 15, 16] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear. INN facilitates three major tasks in computational science and engineering: 1) trains on both experimental and simulation data with very high accuracy and minimal memory footprint, 2) solves high-dimensional parameterized partial differential equations (PDEs), and 3) calibrates uncertain parameters rapidly and accurately.
@@
Figure 2: Benchmark comparison of the single-scale finite difference method (FDM) solver [18], the variational multiscale FEM solver [11], and the INN solver. The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed. A 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM. Detailed problem definition and explanation can be found in Section 3. All solvers are graphics processing units (GPU) optimized with the JAX library [19].
2 INN: Unification of Training/Learning, Solving, and Calibrating
Interpolating Neural Network (INN) generalizes numerical methods such as the finite element method (FEM) [20, 21, 22] as a subset of deep neural network (DNN) via graph network operations. Although different groups of researchers have developed the two (FEM and DNN), they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. INN is built upon this similarity and forms the backbone of Engineering Software 2.0 that unifies training, solving, and calibrating, as illustrated in Figure 1.
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of an INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input – as illustrated in the INN box of Figure 1. The conversion from mesh to graph has been widely accepted in the literature [23, 24, 25, 26]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each segment occupies a subspace of the $I$-dimensional input domain. For instance in Figure 1, the Euclidean mesh consists of segments of hexahedrons and triangular pillars.
Each node (e.g., node $j$) has two features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$ that stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [27]. In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in computational mechanics) for each node $(j=1,\dots,J)$ as a hidden state. A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
@@ 2 INN: Unification of Training/Learning, Solving, and Calibrating
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Supplementary Information (SI) Section 1.1 for visual illustrations). It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$. The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$. This can be any interpolation method such as Lagrange polynomials [28, 20], meshfree [29, 30, 31], and splines [32, 33]. The $Q$-step message passing implies that it is also possible to progressively construct nonlinear or more problem-specific interpolation functions, theoretically based on the generalized finite element method [34, 35] and convolution hierarchical deep learning neural networks [7, 15]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 15]. See SI Section 1 for various choices of the message passing operation.
@@
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
@@
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. Here, $\bm{U}\times^{1}_{2}\bm{\mathcal{X}(x)}$ denotes the tensor contraction operation between the second dimension of tensor $\bm{U}$ and the vector (or the first dimension of tensor) $\bm{\mathcal{X}(x)}$, while $\bm{U}\bm{\mathcal{X}(x)}$ is a matrix-vector multiplication. The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
@@
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of the INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [21, 22]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating.
Special case 1: regular mesh with Tucker decomposition
When the input domain is discretized with a regular mesh, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [13]. One of the widely used TD methods is Tucker decomposition [36, 37]. It approximates a high-order tensor with $J_{i}$ nodes in $i$-th dimension as a tensor contraction between dimension-wise matrices and a core tensor $\bm{\mathcal{G}}$, which has the same order of the original tensor but with smaller nodes $M_{i}$ ($<J_{i}$). The $M_{i}$ is often called a "mode" to distinguish it from the original node $J_{i}$.
Consider a three-input ($I=3$) and one output ($L=1$) system and assume the input domain is discretized with $J=J_{1}\times J_{2}\times J_{3}$ nodes, as shown in the special case box of Figure 1. To facilitate tensor notation, the nodal values will be denoted with left/right super/sub scripts, $\prescript{m}{i}{u}^{(j)}_{l}$, where $i\in\mathbb{N}^{I}$ is the input index, $m\in\mathbb{N}^{M_{i}}$ is the mode index, $l\in\mathbb{N}^{L}$ is the output index, and $j\in\mathbb{N}^{J_{i}}$ is the nodal index.
where the core tensor $\bm{\mathcal{G}}\in\mathbb{R}^{M_{1}\times M_{2}\times M_{3}}$ is a trainable full matrix typically smaller than the original tensor that compresses data. The $\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{M_{i}\times 1}$ is one-dimensional (1D) interpolated output of $i$-th input dimension over $M_{i}$ modes represented as:
$\mathcal{J}\prescript{}{i}{\bm{u}}(x_{i})=\begin{bmatrix}\prescript{1}{i}{u}^{(1)}&\prescript{1}{i}{u}^{(2)}&\cdots&\prescript{1}{i}{u}^{(J_{i})}\\ \prescript{2}{i}{u}^{(1)}&\prescript{2}{i}{u}^{(2)}&\cdots&\prescript{2}{i}{u}^{(J_{i})}\\ \vdots&\vdots&\vdots&\vdots\\ \prescript{M_{i}}{i}{u}^{(1)}&\prescript{M_{i}}{i}{u}^{(2)}&\cdots&\prescript{M_{i}}{i}{u}^{(J_{i})}\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(x_{i})\\ N^{(2)}(x_{i})\\ \vdots\\ N^{(J_{i})}(x_{i})\end{bmatrix}=\prescript{}{i}{\bm{U}}\times^{1}_{2}\prescript{}{i}{\bm{\mathcal{X}(x)}}=\prescript{}{i}{\bm{U}}\prescript{}{i}{\bm{\mathcal{X}(x)}},$ (5)
where $\prescript{}{i}{\bm{U}}\in\mathbb{R}^{M_{i}\times J_{i}}$ and $\prescript{}{i}{\bm{\mathcal{X}(x)}}\in\mathbb{R}^{J_{i}}$. As illustrated in Figure 1’s special case box, the message passing (blue arrow) only happens in the axial directions, yielding 1D interpolation functions: $N^{(J_{i})}(x_{i})$.
When there are more than one outputs $(L>1)$, the interpolated field becomes a vector of $L$ elements:
Special case 2: regular mesh with CP decomposition
where $\mathcal{J}\prescript{m}{i}{\bm{u}}(x_{i})\in\mathbb{R}^{L}$ and $\odot$ represents element-wise multiplication. In CP decomposition, the core tensor is no longer trainable, thus the total trainable parameter becomes $ML\sum^{I}_{i}{J_{i}}$.
It is important to note that the interpolation function $N^{(j)}(x_{i})$ in Eq.5 is one-dimensional. In other words, both Tucker and CP decomposition replace a high-dimensional interpolation with one-dimensional interpolations that dramatically reduce the number of trainable parameters (or DoFs). As given in Table 1, the number of trainable parameters of the full interpolation scales exponentially with the discretization $J_{i}$, whereas the Tucker and the CP decomposition scale linearly with $J_{i}$. Considering the trainable parameters of multi-layer perception (MLP) scales quadratically with the number of hidden neurons, INNs with tensor decomposition (TD) may dramatically reduce the model complexity and computing requirements.
Full interpolation Tucker decomposition CP decomposition
INN trainer
A general INN forward propagation is provided in Eq.3 as a tensor contraction. The INN trainer optimizes nodal values $\bm{U}=\{\bm{u}^{(j)}\}_{j=1,\cdots,J}$ (and if needed, nodal coordinates ($\bm{X}=\{\bm{x}^{(j)}\}_{j=1,\cdots,J}$)) under a given loss function and training data. Consider a regression problem with $K$ labeled data: $(\bm{x}^{*}_{k},\bm{u}^{*}_{k})$, $k=1,\cdots,K$; and $\bm{x}^{*}_{k}\in\mathbb{R}^{I},\bm{u}^{*}_{k}\in\mathbb{R}^{L}$. The superscript * denotes data. A mean squared error (MSE) loss function for this regression problem is defined as:
Finally, an optimization is formulated as:
$\underset{\bm{U},\bm{X}}{\text{minimize}}\ loss(\bm{U},\bm{X}).$ (9)
Section 3.1 exhibits the application of INN trainer to metal additive manufacturing while other benchmarks are elucidated in SI Section 2.
INN solver
An INN solver generalizes classical numerical methods such as FEM and meshfree, as well as the model order reduction methods such as proper generalized decomposition (PGD) [41] and tensor decomposition (TD) [14, 42]. Here, we introduce a formulation with CP decomposition to solve the generalized space $(\bm{x})$ time $(t)$ parameter $(\bm{\theta})$ (S-T-P) PDE, whose computational cost is prohibitively high for most numerical methods and machine learning approaches [41].
Consider a parameterized space-time PDE:
$\mathcal{L}\bm{u}(\bm{x},t,\bm{\theta})=0,$ (10)
[…]
where $\delta\bm{u}$ is the weight function and can be defined using Galerkin, Petrov-Galerkin, collocation or quadratic formulation. Due to the Kronecker delta property of INN interpolation functions, Dirichlet boundary conditions and initial conditions can be strongly imposed. As a result, only the weighted sum residual is considered in Eq. 11 as a loss function. This distinguishes INNs from most data-driven PDE solvers that weakly impose those conditions [46]. See the SI Section 4 for detailed derivations of the loss function and solution scheme. INN solvers applied to the metal additive manufacturing and the linear elastic solid mechanics problem are introduced in Section 3.3 and SI Section 4, respectively.
INN calibrator
[…]
3 Scientific and Engineering applications: Metal Additive Manufacturing

### v5

# Interpolating neural network: A novel unification of machine learning and interpolation theory

@@ Abstract
Artificial intelligence (AI) has revolutionized software development, shifting from task-specific codes (Software 1.0) to neural network-based approaches (Software 2.0). However, applying this transition in engineering software presents challenges, including low surrogate model accuracy, the curse of dimensionality in inverse design, and rising complexity in physical simulations. We introduce an interpolating neural network (INN), grounded in numerical methods, to realize Engineering Software 2.0 by advancing data training, partial differential equation solving, and parameter calibration. INN offers orders of magnitude fewer trainable/solvable parameters for comparable model accuracy than traditional multi-layer perceptron (MLP) or physics-informed neural networks (PINN). Demonstrated in metal additive manufacturing, INN rapidly constructs an accurate surrogate model of Laser Powder Bed Fusion (L-PBF) heat transfer simulation, achieving sub-10-micrometer resolution for a 10 mm path in under 15 minutes on a single GPU. This makes a transformative step forward across all domains essential to engineering software.
@@ 1 Introduction
The evolution of software programming methodologies has transitioned from reliance on rigidly hard-coded instructions, governed by explicitly defined human rules, to the adoption of neural network-based algorithms. The transition is coined as "from Software 1.0 to Software 2.0" [1]. The shift towards Software 2.0 partially resolved the issue of labor-intensive programming in Software 1.0 and has significantly advanced the domain of large language models (LLMs) and other foundational models [2, 3]. However, the application of these technologies in the fields of engineering and science presents unique challenges. One such challenge is the relatively low generalization accuracy of machine learning (ML)-based partial differential equation (PDE) solvers compared to traditional numerical methods [4, 5, 6, 7, 8]. After in-depth research was conducted on those methods, an increasing number of researchers began to question their effectiveness [8, 9]. Another challenge is the curse of dimensionality that arises when solving inverse design problems. As the design parameter space grows, computation requirements easily become prohibitive. Furthermore, the complexity of physical simulations continues to increase. For instance, additive manufacturing (AM) or integrated circuit (IC) simulations necessitate extremely fine resolution with high accuracy and multi-physics/multi-scale analysis [10, 11, 12].
@@
In this article, we propose the concept of Engineering Software 2.0, which aims to advance the current generation of numerical and data-driven engineering software into a new paradigm.
@@
“Engineering Software 2.0 is an end-to-end software system that unifies and advances data training, PDE solving, and parameter calibration in science/engineering problems."
We introduce interpolating neural networks (INNs) to achieve the new paradigm of Engineering Software 2.0. INNs generalize numerical methods such as the finite element method (FEM) [13, 14, 15] as a subset of deep neural network (DNN) or graph neural network (GNN). Although different groups of researchers have developed the two (FEM and neural networks), they share one thing in common; the goal is to find a function. The former finds a solution field of a differential equation while the latter finds an input-output relationship. Although it is possible to formulate INN from a DNN, we adopt GNN architecture to explain INN as a special case of GNN with a set of rules (i.e., interpolation).
GNNs provide a general formulation to process information available in a graph, i.e. a cloud of interconnected points or nodes [16, 17, 18]. Each node receives information from its connected neighbors that is then processed by a feedforward neural network to make a new prediction at that node. However, their network architectures are often overly complex, making convergence difficult for highly nonlinear problems. In Science and Engineering, most computer simulations also occur in interconnected point clouds (or nodes), but each interconnected node is associated with simpler interpolatory functions that are fast to differentiate and easy to integrate, leading to methods that obey established error bounds and convergence rates. This provides confidence in the predictions made but also imposes a need for relatively fine discretizations (dense point clouds) because each point is related to its connected neighbors by a simple function (e.g., low order polynomials or splines). INNs bridge the gap between the extreme generality of GNNs and the stringent restrictions imposed by numerical methods by constructing interpolatory GNNs that can simplify to known numerical methods or complexify to known GNN architectures.
@@
INNs consist of three processes: 1) discretize an input domain into non-overlapping segments whose bounds are denoted by interpolation nodes, 2) construct a graph with the interpolation nodes and formulate the message passing operation as a form of interpolation functions, and 3) optimize the values and coordinates of the interpolation nodes for a given loss function. Figure 1 describes the approach. When the input domain is discretized with a regular mesh (see the special case in Figure 1), INNs can leverage tensor decomposition (TD) [19, 20, 21, 22] to convert the growth of the computational cost of a high-dimensional problem from exponential to linear. INNs facilitate three major tasks in computational science and engineering: 1) data training, 2) PDE solving, and 3) parameter calibration, by reducing trainable parameters, training time, and memory/storage requirements without compromising model accuracy.
@@
Figure 2: Illustration of a 1D input domain discretized with 5 segments and 6 nodes.
2 Results
Consider a regression problem that relates $I$ inputs and $L$ outputs. The first step of an INN is to discretize the input domain in the $I$-dimensional Euclidean space into a mesh, which can be as general as an unstructured irregular mesh or as specific as a structured regular mesh. In any case, a mesh in the Euclidean space can be readily represented as a graph – the most general form of a discretized input – as illustrated in the INN box of Figure 1. The conversion from mesh to graph has been widely accepted in the literature [25, 26, 27, 28]. Suppose there are $J$ nodes (or vertices) and $E$ edges in the graph, where the input domain is discretized with $C$ non-overlapping segments. Each segment occupies a subspace of the $I$-dimensional input domain.
Each node (e.g., node $j$) has features: nodal coordinates $(\bm{x}^{(j)}\in\mathbb{R}^{I})$ and values $(\bm{u}^{(j)}\in\mathbb{R}^{L})$. Note that the superscript with parentheses refers to the graph node index. An edge that links node $j$ and $k$ has a feature $\bm{e}^{(j,k)}$ that stores indices of segments connected to the edge. For example, when $I=3,L=2$, node $10$ in Figure 1 stores $\bm{x}^{(10)}=(x^{(10)}_{1},x^{(10)}_{2},x^{(10)}_{3})$ and $\bm{u}^{(10)}=(u^{(10)}_{1},u^{(10)}_{2})$. The edge connecting nodes $30$ and $12$ stores $\bm{e}^{(30,12)}=\{5,7\}$.
A typical message passing in a graph neural network (GNN) returns a hidden state for each graph node which is mostly a tensor (including matrix and vector) [16]. In contrast, the INN message passing returns a function (i.e., interpolation function $N^{(j)}(\bm{x})$, or often called shape function in FEM) for each node $(j=1,\dots,J)$ as a hidden state. A general $Q$-step message passing of an INN, $\mathcal{M}^{[q]}(*)$, can be expressed as:
@@ 2 Results
where $N^{[q](j)}(\bm{x})$ is the interpolation function at node $j$ after $q$-th message passing and $\mathcal{N}^{s}(j)$ is a set of neighboring nodes of the center node $j$ with $s$ connections (i.e., $s$-hops, see Figure 2 and Supplementary Information (SI) Section 1.1 for visual illustrations). It is worth noting that the interpolation functions satisfy the Kronecker delta property, i.e., $N^{[q](j)}(\bm{x}^{(k)})=\delta_{jk}$. The operation $\mathcal{M}^{[q]}(*)$ constructs an interpolation function $N^{[q](j)}(\bm{x})$ for a graph node $\bm{x}^{(j)}$ using neighboring nodal coordinates and edge information: ${\bm{x}^{(k)},\bm{e}^{(j,k)}}$, and the interpolation function of the previous message passing: $N^{[q-1](j)}(\bm{x})$.
When there is only one message passing ($Q=1$) with $s=1$ hop, INNs degenerate to the standard FEM linear shape function (or piecewise linear interpolation). As illustrated in Figure 2 (left), FEM with linear elements is the most localized approximation of an INN. On the other hand, one can progressively enlarge the support domain and the approximation capability of an interpolation function by adding message passings with a higher hop. For instance, the second message passing with $s=2$ hop constructs compact-supported interpolation functions with higher nonlinearity. This is theoretically equivalent to the generalized finite element method (GFEM) [29] and convolution hierarchical deep learning neural networks (C-HiDeNN) [7, 21]. The message passing with a max-hop (Figure 2, right) can even make it a non-local approximation where support of the interpolation functions occupy the global domain, mimicking meshfree methods that exhibit superconvergence (i.e., faster convergence rate than the complete order of polynomial basis) [30, 31, 32]. Depending on the choice of interpolation technique, other hyper-parameters can be involved in $\mathcal{M}^{[q]}(*)$ such as the activation (or basis) function and dilation parameter [7, 21]. See SI Section 1 for various choices of the message passing operation.
Regardless of the choice of interpolation technique for the message passing, the global degrees of freedom (i.e., number of nodes) remain constant. This distinguishes INNs from higher-order FEM where each segment is enriched with additional nodes. INNs, on the other hand, adapt nodal connectivity through the adjacency matrix and basis functions to reproduce almost any interpolation technique available in numerical methods. The compact-supported interpolation functions enable INNs to be optimized locally, making INNs distinguishable from MLPs. Since an MLP is a global approximator, INNs with compact-supported interpolation functions 1) converge faster than MLPs with the same number of trainable parameters (see Section 2.1) and 2) facilitate training on a sparse dataset (see Section 3.2), although it might end up with overfitting.
@@
$\mathcal{J}\bm{u}(\bm{x})=\begin{bmatrix}|&|&|&|\\ \bm{u}^{(1)}&\bm{u}^{(2)}&\cdots&\bm{u}^{(J)}\\ |&|&|&|\end{bmatrix}\cdot\begin{bmatrix}N^{(1)}(\bm{x})\\ N^{(2)}(\bm{x})\\ \vdots\\ N^{(J)}(\bm{x})\end{bmatrix}=\bm{U}\bm{\mathcal{X}(x)},$ (3)
@@
where $\bm{u}^{(j)}\in\mathbb{R}^{L},\bm{U}\in\mathbb{R}^{L\times J},\bm{\mathcal{X}(x)}\in\mathbb{R}^{J}$. The matrix $\bm{U}$ is a stack of nodal values $\bm{u}^{(j)}$, while $\bm{\mathcal{X}(x)}$ is a vectorized function of $\bm{x}$ that is parameterized with nodal coordinates $\bm{x}^{(j)}$ in the message passing operation.
@@
The graph node features (i.e., coordinate ($\bm{x}^{(j)}$) and value ($\bm{u}^{(j)}$)) are trainable parameters of the INN. If the nodal coordinates ($\bm{x}^{(j)}$) are fixed, one can find nodal values ($\bm{u}^{(j)}$) without changing the discretization of the input domain. If the nodal coordinates are also updated, the optimization will adjust the domain discretization similar to r-adaptivity in FEM [14, 15]. Once the forward propagation is defined, the loss function is chosen based on the problem type: training, solving, or calibrating (see Section 5 for details).
When the input domain is discretized with a regular mesh, as illustrated in the special case of Figure 1, we can significantly reduce the trainable parameters (or degrees of freedom, DoFs) by leveraging Tensor Decomposition (TD) [19]. Section 5 introduces the two widely accepted TD methods: Tucker decomposition [33, 34] and CANDECOMP/PARAFAC (CP) decomposition [35, 36, 37].
Figure 3: Benchmarks for the INN solver (a) and the INN trainer (b). In (a), a 3D-space, time, and 4D-parameter heat transfer equation is solved to model a 10 mm single-track laser powder bed fusion (L-PBF) metal AM. Detailed problem definition and explanation can be found in Section 3. We compare the single-scale finite difference method (FDM) solver [24], the variational multiscale FEM solver [11], and the INN solver with CP decomposition and Q=2. The data points for the first two methods with dashed marker edges are estimated, while that of INN with a solid marker edge is computed. In (b), a standard data training (regression) problem is studied. Each trainer stops when the training loss hits the stopping criteria (MSE:4e-4). We set the same optimization condition: ADAM optimizer; learning rate: 1e-3; batch size: 128. The number of neurons per hidden layer is denoted for each MLP data point while the number of segments and modes are denoted for each INN data point. INN with CP decomposition and Q=2 is used. In (c), a 1D Poisson’s equation is solved with PINN and INN. PINN is made of a 1-layer MLP with a varying number of neurons. Randomly selected 10k collocation points are used to compute the PDE loss. With a batch size of 128, 10k epochs are trained using an ADAM optimizer with a learning rate of 1e-1. FEM with linear elements is equivalent to an INN with Q=1, s=1. C-HiDeNNs with s=3,P=3 and with s=5,P=5 are equivalent to INNs with Q=2, s=3, P=3, and with Q=2, s=5, P=5, respectively. All solvers and trainers are graphics processing units (GPU) optimized with the JAX library [38].
2.1 Benchmarking INN trainer
To highlight INN’s advantages in speed (training epoch) and storage (number of parameters) over traditional MLP, we introduce a benchmark problem: a 10-input 5-output physical function [39], (see SI Section 2.4 for details). Using this deterministic equation, we randomly generate 100,000 data using a Latin hypercube sampling. The dataset is split into 80% and 20% for train and test, respectively. MLPs with two and three hidden layers and with a sigmoid activation are tested while INNs with CP decomposition, two (Q=2) message passing, and $s=2,P=2$ polynomial activation are adopted. To investigate the convergence behavior depending on the number of parameters, we set the stopping criteria as training MSE: 4e-4 and count the epoch at convergence. Since the physical function is deterministic (no noise) and we drew a sufficiently large number of data, no considerable overfitting was observed.
Figure 3(b) reveals that given the same number of trainable parameters, INNs converge significantly faster than MLPs. For instance, the INN with 2,250 parameters converges at the 34th epoch while the MLP with 2,285 parameters converges at the 217th epoch. MLPs with more parameters tend to stop at earlier epochs, however, even the largest MLP with 21,805 parameters converges at the 45th epoch while that of INN with only 8,250 parameters converges at the 13th epoch. This benchmark demonstrates that the INN trainer is lightweight and fast-converging compared to traditional MLPs. Other benchmarks of the INN trainer are elucidated in SI Section 2.
2.2 Benchmarking INN solver
INNs can achieve a proven convergence rate when solving a PDE. In numerical analysis, the convergence rate is the decaying rate of an error measurement as the mesh refines. This benchmark solves a 1D Poisson’s equation defined in Section 5.6 with the $H^{1}$ norm error estimator. It is mathematically proven that a numerical solution of FEM or any interpolation-based solution method exhibits a convergence rate of $P$, which is the complete order of polynomial basis used in the interpolation [21].
Figure 3(c) is the $H^{1}$ error vs. parameters (i.e., degrees of freedom) plot where the convergence rates are denoted as the slope of the log-log plot. Depending on the message passing operation, INNs degenerate to linear FEM ($P=1$) or C-HiDeNNs (or GFEM) with different order of polynomials: $P=3,P=5$. However, PINNs constructed with MLP do not reveal a convergent behavior. Only a few theoretical works have proved the convergence rate of PINNs under specific conditions [40]. A PDE solver needs to have a stable and fast convergence rate because it guides an engineer in choosing the mesh resolution and other hyperparameters for achieving the desired level of accuracy. We also demonstrate the convergent behavior of INN solvers with and without CP decomposition for a 3D linear elasticity equation in SI Sections 4.5 and 4.6.
3 Discussion